OpenStudy (anonymous):

Bridget has a kitchen canister, in the shape of a cylinder, which holds flour. If the canister is 24 centimeters wide and 26 centimeters tall, what is its approximate surface area in terms of pi? 624π square centimeters 912π square centimeters 1,248π square centimeters 1,296π square centimeters

OpenStudy (saifoo.khan):

Apply this formula. http://0.tqn.com/d/math/1/5/F/F/Cylinderr.gif

OpenStudy (campbell_st):

surface area of a closed cylinder is \[SA = 2\pi r^2 + 2\pi r h\] have the wide to find the radius of the cylinder... and you need to decide if the cylinder is closed... of open at one end. If open at one end is \[SA = \pi r^2 + 2\pi rh\]

OpenStudy (anonymous):

I tried to figure this out, but I got 1959.39

OpenStudy (anonymous):

I ment .36

OpenStudy (campbell_st):

just do the calculation with out pi... as the answers are written in exact form So do the calculations with out pi... and then you should get a number and add a pi at the end

OpenStudy (anonymous):

i dont get this. i have the same question but i dont get exactly how to do it ffrom what yall explained it. can you solve this problem step by step so i can see how to exactly do this